OpenStudy (anonymous):

1.write an expression to represent the time you still have left to finish an exam if the whole exam takes "t" hours and you have been working on the exam for 25 minutes. a. give the answer in minutes b.give the answer in hours anyone?? needd help >..<

OpenStudy (anonymous):

in minutes it would be t/25=t amount of minutes. ex. if the test took 5 hors the first thing you would do is 60x5(which equals 300 btw) then you would just do 300/25 which equals 12 Hours are a bit trickier. I think that because there are 60 mins in one hour and thirty mins in a half hour, it would be either 60/25 or 30/25. 30/25=1.2 and 60/25 is double the ammount (2.4)

OpenStudy (anonymous):

thats what i think though

OpenStudy (anonymous):

i hope im right! :D

OpenStudy (anonymous):

ty for the help :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ohh oohh hmm so 1.2 is the minutes and 2.4 is the hour???

OpenStudy (anonymous):

not really, because we dont know what "t" is, so its an open equation ( t/25=t amount of minutes.) however, i dont know if they count half hours and/or whole hours, so i put in the half hour and the whole hour that way i can count both

OpenStudy (anonymous):

you understand what i mean right?

OpenStudy (anonymous):

ok ok :)

OpenStudy (anonymous):

alright cool! again i hope i get the correct answer, because im not one for giving wrong answers, lol